OpenStudy (anonymous):

Evaluate the expression for a = -1 and b=5. (ab)2

OpenStudy (anonymous):

25

OpenStudy (anonymous):

(-1*5)^2=25

OpenStudy (mathstudent55):

(ab)^2 = ( ( -1) * 5 )^2 = (-5)^2 = 25